# PIR: Remote Sensing Image-Text Retrieval with Prior Instruction Representation Learning

### v1

@@ III-E2 Affiliation Loss
III-F Training Paradigm
For PIR-ITR with soft belief strategy, we train on the RSICD dataset [8] or RSITMD dataset [9] as shown in Algorithm 1. We first obtain positive remote sensing image-text pairs from the remote sensing image-text datasets . For vision and text encoding, the image and text are encoded separately to obtain the corresponding visual features , instruction embeddings and textual features . For vision instruction representation, we compute the belief matrix of visual features and instruction embeddings and calculate the rank of element in batch training. For each element of dimension in , we get rank-weighted visual features . Then we can use Spatial-PAE and Temporal-PAE to get and respectively. Notably, the clustering centers for each scene category of each batch need to be computed before calculating affiliation loss. Finally we compute the contrastive loss and affiliation loss to update the model parameters that require gradient updating.
The PIR-ITR directly performs the PIR learning paradigm after visual and text encoder processing for the above closed-domain retrieval. However, this complex and large structure is not easy to train on large-scale data. For PIR-CLIP, we first pre-train on RS5M dataset [29] using CLIP with contrastive loss, then fine-tune visual representation from RSITR datasets, the fine-tuning procedure of PIR-CLIP as shown in Algorithm 2. The specific fine-tuning process is similar to that of PIR-ITR described above, with the difference that VIR is modified with the base ViT and ResNet models. LCA is replaced with a language Transformer trained on a large-scale remote sensing corpus. Compared to PIR-ITR, PIR-CLIP is a modification of the ViT structure, using convolution to accomplish patching, followed by a soft belief strategy, and then a Transformer encoder to interact between vision features and instruction features. The language encoder trained with the help of large-scale image-text pairs can directly encode the final text embedding without further post-processing by LCA.
IV EXPERIMENTS
IV-A Datasets
In our experiments, we use two benchmark datasets RISCD [8] and RSITMD [9], which are commonly used in RSITR [52, 9, 23, 24].
IV-A1 RSICD Dataset
IV-A2 RSITMD Dataset
@@ IV-B Metrics
IV-C Implementation Details
For closed-domain retrieval, Swin Transformer is used as the vision encoder, pre-trained on ImageNet [70] using Swin-T (tiny version) 222https://github.com/microsoft/Swin-Transformer, and BERT is used as the text encoder with pre-trained parameters from BERT-B (base version) provided by the official source 333https://github.com/google-research/bert. The output dimensions for both vision and text encodings are set to 768 and then linearly mapped to 512 dimensions. Both self-attention and cross-attention have a head size of 8 and a dropout rate of 0.2, with different numbers of repeat units (2 for Spatial-PAE and 3 for Temporal-PAE). ResNet-50 pre-trained on AID dataset [61] is used as the instruction encoder, with the second-to-last layer as an output of size 1024, linearly mapped to 512 dimensions. The contrastive and attribution losses have a temperature coefficient of 0.07 and a center factor of 1. Training is performed in parallel on 2 NVIDIA RTX A6000 GPUs, with a batch size of 128, using AdamW [71] optimizer with a learning rate of and weight decay of 0.01. VIR filter size is set to 40 both on RSITMD and RSICD.
For open-domain retrieval, ResNet-50 + ViT-B and Transformer are used as visual and textual backbones to align vanilla CLIP. We train a modified ResNet-50 on the AID dataset [61] and freeze the parameters before the attention pool layer during CLIP tuning. The pre-training epoch is set to 20, the batch size to 256, the fine-tuning is set to 10, and the batch size to 512, and the training samples go through 1,000,000 steps per epoch. The pre-training and fine-tuning of PIR-CLIP is performed on a cluster of 4 NVIDIA A100 GPUs with learning rate and using automatic mixed-precision. Other settings are the same as for PIR-ITR. Model codes and checkpoints are open-sourced at https://github.com/jaychempan/PIR-CLIP.
@@ IV-C Implementation Details
IV-D Performance Comparisons
In this section, we compare the proposed PIR-ITR with the state-of-the-art methods on RSICD and RSITMD. Image-query-Text (I2T) and Text-query-Image (T2I) retrieval results are counted for the experiments. Experiments are repeated three times to ensure reliability, and then the average results are taken as the final results.
IV-D1 State-of-the-art methods
@@
Closed-domain methods can be divided into three groups: traditional image-text retrieval methods, [72], SCAN [48], CAMP [52], and CAMERA [73], remote sensing image-text retrieval methods, LW-MCR [52], AMFMN [9], GaLR [23], KCR [22], SWAN [24], HVSA [74], and DOVE [75], and Transformer-based backbone methods, , [72] and .
Most of the above methods are based on different vision and text feature extraction backbones. We use the provided source code for implementation for traditional image-text retrieval methods. For SCAN, CAMP, and CAMERA, RoI Trans [77] with ResNet-50 as backbone 444https://github.com/open-mmlab/mmrotate pre-trained on DOTA [78] is used to detect significant objects. For remote sensing image-text retrieval methods, we pick the best results from the original paper to make the experiments as fair as possible since the experiments are difficult to be reproduced on RSICD and RSITMD. To be as fair as possible, we add Transformer-based backbone methods 555In closed-domain methods, most of them use BERT for text feature extraction, different from open-domain methods that use vanilla Transformer for the purpose of establishing a foundation model. compared with PIR-ITR, a further improvement of the method. represents different vision and text backbones and TEL denotes the Transformer encoder layer. We also marked the corresponding loss function on the method (TL: triplet loss, CL: contrastive loss). “PIR-ITR with hard” and “PIR-ITR with soft” is PIR-ITR with hard and soft belief strategy.
@@
IV-D2 Quantitative comparison of Closed-domain methods
@@
IV-D3 Quantitative comparison of Open-domain methods
@@ IV-D4 Visualization of results
IV-E Ablation Studies
In this section, we design twelve sets of experiments (the first eight is for PIR-ITR, else for PIR-CLIP) to demonstrate the effectiveness of the proposed three modules VIR, LCR and affiliation loss (), as shown in Table III. The baseline group represents the removal of the above three modules as the blank control group. CLIP-baseline is PIR-CLIP removes VIR and attribution loss; other conditions remain the same.
IV-E1 Effects of Vision Instruction Representation
From the experimental results of +VIR, we find that using VIR alone leads to a small increase in compared to the experimental results of baseline, with a significant increase in T2I . From the experimental results of +VIR+LCA, we find that VIR increases from 36.95 to 37.25 compared to the experimental results of +LCA, where VIR increases more for I2T and T2I . From the experimental results of +VIR+LCA+, is improved by 2.6% compared to the experimental results of +LCA+, where I2T increases from to and T2I increases from to . For PIR-CLIP, we find that VIR increases from 45.35 to 48.02 compared with the experimental results of +VIR and CLIP-baseline, and increases from 48.16 to 52.68 compared with the experimental results of + and +VIR +. With the above results, we find that VIR can significantly improve vision representation and enhance the ability of Text-query-Image.
@@ IV-E1 Effects of Vision Instruction Representation
IV-E2 Effects of Language Cycle Attention
@@
From the experimental results of +LCA, we find that LCA can respectively improve 24.6% and 35.8% on I2T and T2I compared to the experimental results of baseline. From the experimental results of +VIR+LCA, we can observe that a 1.2% improvement in compared to the experimental results of +VIR. From the experimental results of +VIR+LCA+, LCA results in a 1.4% improvement in compared to the experimental results of +VIR+, where I2T is improved from 16.81 to 18.14 and T2I is improved from 11.59 to 12.17. Comparing the experimental group +VIR with +VIR+LCA and the experimental group + with +LCA+, we find that the effect of LCA on the Image-query-Text and Text-query-Image task is positive. With the above results, we find that LCA balances with vision representation to improve retrieval performance.
IV-E3 Effects of Affiliation Loss
@@
From the experimental results of +, we find that alone can significantly improve I2T compared to the experimental results of baseline, which is improved from 16.15 to 19.25. From the experimental results of +VIR+LCA+, it can be observed affiliation loss could improve by 2.7% compared to the experimental results of +VIR+LCA, where I2T is improved from 17.04 to 18.14. For PIR-CLIP, we find that increases from 45.35 to 48.16 compared with the experimental results of + and CLIP-baseline, and increases from 48.02 to 52.68 compared with the experimental results of +VIR and +VIR +. From the above results, we can conclude that affiliation loss can significantly improve vision and text representations, which can help VIR and LCA to perform long-range dependency modeling.
IV-F Parameter Evaluation
There are two important parameters in our proposed modules: the filter size in VIR and the center scale of affiliation loss. This section focuses on experiments to select the appropriate parameters and verify further the roles played in the retrieval process.
IV-F1 Filter Size in VIR
@@
For PIR-ITR with hard belief strategy, the visual features extracted based on Transformer contain much semantic noise, and some of the key features are more representative of the visual representation, so we use VIR for unbiased correction of the visual representation. The maximum number of Vision Transformer output features is 50, and we control the number of key features by adjusting the filter size. Table IV shows the bidirectional retrieval results on the RSITMD with different filter sizes (see Equation 9). We can find that when =10, I2T is the largest, reaching 54.65; when =20, reaches 38.00, and T2I is the largest, reaching 13.50; when =30, I2T is the largest, but are relatively low compared to others; when =40, is the largest, reaching 38.24, and the key feature is more representative of visual features at this time; when =50, is 37.48 smaller than =10, 20, 40 indicating that some key features are susceptible to deactivation by semantic noise. For PIR-ITR with soft belief strategy, it can further enhance the image retrieval and text retrieval performance with reaching up to 39.25. From the above experimental results, it can be concluded that semantic noise contained in remote-sensing images deactivates key features, and the retrieval performance can be effectively improved by reducing the dependence on partial features.
IV-F2 Center Scale of Affiliation Loss
@@
IV-G Instruction Strategy Analysis
To explore the impact of different instruction strategies on retrieval performance, we use different types of models as instruction encoders on PIR-ITR 666PIR-CLIP is designed to further explore the action of PIR in the foundation model, thus these experiments are based on PIR-ITR. with hard belief strategy, where each model is pre-trained on the AID [61], RESISC45 dataset [80] and ImageNet [81] datasets for the classification task. These are divided into two categories, the CNN-based encoders, including VGG16 [82], VGG-19 [82], ResNet-50 [60], and ResNet-101 [60], and the Transformer-based encoders, including ViT [35] and Swin-T [63]. Table V shows the experimental results of the instruction encoders on the RSITMD dataset. The experimental results show that using ResNet-50 pre-trained on the AID dataset and RESISC45 dataset as the instruction encoder is more effective, performing the best at 38.24 and 38.28. We find no significant increase on retrieval performance by looking at the pre-trained ViT and Swin-T on the AID dataset as instruction encoders. By comparing VGG-16 and ResNet-50 pre-trained on ImageNet, we find that the pre-training based on remote sensing scene classification has a more obvious instructive role in guiding the vision representation. The above experimental results show that using prior knowledge based on remote sensing scene recognition can enhance vision representation and, thus, retrieval performance.
@@ IV-G Instruction Strategy Analysis
V CONCLUSION
This paper proposes the PIR framework, a new paradigm that leverages prior knowledge from the remote sensing data domain for adaptive learning of vision and text representations. We first propose the PIR-ITR for remote sensing image-text retrieval. Notably, Spatial-PAE and Temporal-PAE are applied to perform long-range dependency modelling to enhance key feature representation. In vision representation, VIR based on Spatial-PAE utilizes prior knowledge from remote sensing scene recognition to instruct unbiased vision representation and reduce semantic noise’s impact. In text representation, LCA based on Temporal-PAE activates the current time step based on the previous time step to enhance semantic representation capability. Finally, a cluster-wise attribution loss is used as an inter-class constraint to reduce semantic confusion zones in the common subspace. Building on this foundation, we propose PIR-CLIP to address semantic noise in remote sensing vision-language representations and enhance open-domain retrieval performance further. Comprehensive experiments are designed to validate the superiority and effectiveness of our approach on two benchmark datasets, RSICD and RSITMD.
In the future, a unified multimodal remote sensing foundation model that is more specific to understanding complex multimodal data from remote sensing should be proposed.

### v2

@@ IV-D2 Affiliation Loss
V EXPERIMENTS
V-A Datasets
V-A1 RSICD Dataset
V-A2 RSITMD Dataset
@@ V-B Metrics
V-C Implementation Details
For closed-domain retrieval, Swin Transformer is used as the vision encoder, pre-trained on ImageNet [58] using Swin-T (tiny version) 222https://github.com/microsoft/Swin-Transformer, and BERT is used as the text encoder with pre-trained parameters from BERT-B (base version) provided by the official source 333https://github.com/google-research/bert. The output dimensions for both vision and text encodings are set to 768 and then linearly mapped to 512 dimensions. Both self-attention and cross-attention have a head size of 8 and a dropout rate of 0.2, with different numbers of repeat units (2 for Spatial-PAE and 3 for Temporal-PAE). ResNet-50 pre-trained on AID dataset [52] is used as the instruction encoder, with the second-to-last layer as an output of size 1024, linearly mapped to 512 dimensions. The contrastive and attribution losses have a temperature coefficient of 0.07 and a center factor of 1. For open-domain retrieval, ResNet-50 + ViT-B and Transformer are used as visual and textual backbones to align vanilla CLIP. We train a modified ResNet-50 on the AID dataset [52] and freeze the parameters before the attention pool layer during CLIP tuning. The pre-training epoch is set to 20, the batch size to 256, the fine-tuning is set to 10, and the batch size to 512, and the training samples go through 1,000,000 steps per epoch. Other settings are the same as for PIR-ITR. Model codes and checkpoints are open-sourced at https://github.com/jaychempan/PIR-CLIP.
@@ V-C Implementation Details
V-D Performance Comparisons
V-D1 State-of-the-art methods
@@
Closed-domain methods can be divided into three groups: traditional image-text retrieval methods, [43], SCAN [35], CAMP [39], and CAMERA [44], remote sensing image-text retrieval methods, LW-MCR [39], AMFMN [21], GaLR [11], KCR [10], SWAN [12], HVSA [59], and DOVE [54], and Transformer-based backbone methods, , [43] and . For SCAN, CAMP, and CAMERA, RoI Trans [61] with ResNet-50 as backbone 444https://github.com/open-mmlab/mmrotate pre-trained on DOTA [62] is used to detect significant objects. For remote sensing image-text retrieval methods, we pick the best results from the original paper to make the experiments as fair as possible since the experiments are difficult to be reproduced on RSICD and RSITMD. To be as fair as possible, we add Transformer-based backbone methods 555In closed-domain methods, most of them use BERT for text feature extraction, different from open-domain methods that use vanilla Transformer for the purpose of establishing a foundation model. compared with PIR-ITR, a further improvement of the method. represents different vision and text backbones and TEL denotes the Transformer encoder layer. We also marked the corresponding loss function on the method (TL: triplet loss, CL: contrastive loss). “PIR-ITR with hard” and “PIR-ITR with soft” is PIR-ITR with hard and soft belief strategy.
@@
V-D2 Quantitative comparison of Closed-domain methods
@@
V-D3 Quantitative comparison of Open-domain methods
@@ V-D4 Visualization of results
V-E Ablation Studies
In this section, we design twelve sets of experiments (the first eight is for PIR-ITR, else for PIR-CLIP) to demonstrate the effectiveness of the proposed three modules Spatial-PAE, Temporal-PAE and affiliation loss (), as shown in Table III. The baseline group represents the removal of the above three modules as the blank control group. CLIP-baseline is PIR-CLIP removes Spatial-PAE and attribution loss; other conditions remain the same.
V-E1 Effects of Vision Instruction Representation
From the experimental results of +S-PAE, we find that using Spatial-PAE alone leads to a small increase in compared to the experimental results of baseline, with a significant increase in T2I . From the experimental results of +S-PAE+T-PAE, we find that Spatial-PAE increases from 36.95 to 37.25 compared to the experimental results of +T-PAE, where Spatial-PAE increases more for I2T and T2I . From the experimental results of +S-PAE+T-PAE+, is improved by 2.6% compared to the experimental results of +T-PAE+, where I2T increases from to and T2I increases from to . For PIR-CLIP, we find that Spatial-PAE increases from 45.35 to 48.02 compared with the experimental results of +S-PAE and CLIP-baseline, and increases from 48.16 to 52.68 compared with the experimental results of + and +S-PAE +. With the above results, we find that Spatial-PAE can significantly improve vision representation and enhance the ability of Text-query-Image.
@@ V-E1 Effects of Vision Instruction Representation
V-E2 Effects of Language Cycle Attention
@@
From the experimental results of +T-PAE, we find that Temporal-PAE can respectively improve 24.6% and 35.8% on I2T and T2I compared to the experimental results of baseline. From the experimental results of +S-PAE+T-PAE, we can observe that a 1.2% improvement in compared to the experimental results of +S-PAE. From the experimental results of +S-PAE+T-PAE+, Temporal-PAE results in a 1.4% improvement in compared to the experimental results of +S-PAE+, where I2T is improved from 16.81 to 18.14 and T2I is improved from 11.59 to 12.17. Comparing the experimental group +S-PAE with +S-PAE+T-PAE and the experimental group + with +T-PAE+, we find that the effect of Temporal-PAE on the Image-query-Text and Text-query-Image task is positive. With the above results, we find that Temporal-PAE balances with vision representation to improve retrieval performance.
V-E3 Effects of Affiliation Loss
@@
From the experimental results of +, we find that alone can significantly improve I2T compared to the experimental results of baseline, which is improved from 16.15 to 19.25. From the experimental results of +S-PAE+T-PAE+, it can be observed affiliation loss could improve by 2.7% compared to the experimental results of +S-PAE+T-PAE, where I2T is improved from 17.04 to 18.14. For PIR-CLIP, we find that increases from 45.35 to 48.16 compared with the experimental results of + and CLIP-baseline, and increases from 48.02 to 52.68 compared with the experimental results of +S-PAE and +S-PAE +. From the above results, we can conclude that affiliation loss can significantly improve vision and text representations, which can help Spatial-PAE and Temporal-PAE to perform long-range dependency modeling.
V-F Parameter Evaluation
V-F1 Filter Size in PIR
@@
For PIR-ITR with hard belief strategy, the visual features extracted based on Transformer contain much semantic noise, and some of the key features are more representative of the visual representation, so we use for unbiased correction of the visual representation. The maximum number of Vision Transformer output features is 50, and we control the number of key features by adjusting the filter size. Table IV shows the bidirectional retrieval results on the RSITMD with different filter sizes (see Equation 4). We can find that when =10, I2T is the largest, reaching 54.65; when =20, reaches 38.00, and T2I is the largest, reaching 13.50; when =30, I2T is the largest, but are relatively low compared to others; when =40, is the largest, reaching 38.24, and the key feature is more representative of visual features at this time; when =50, is 37.48 smaller than =10, 20, 40 indicating that some key features are susceptible to deactivation by semantic noise. For PIR-ITR with soft belief strategy, it can further enhance the image retrieval and text retrieval performance with reaching up to 39.25. From the above experimental results, it can be concluded that semantic noise contained in remote-sensing images deactivates key features, and the retrieval performance can be effectively improved by reducing the dependence on partial features.
V-F2 Center Scale of Affiliation Loss
@@
V-G Instruction Strategy Analysis
To explore the impact of different instruction strategies on retrieval performance, we use different types of models as instruction encoders on PIR-ITR with hard belief strategy, where each model is pre-trained on the AID [52] dataset, RESISC45 dataset [65] and ImageNet [66] datasets for the classification task. These are divided into two categories, the CNN-based encoders, including VGG16 [67], VGG-19 [67], ResNet-50 [51], and ResNet-101 [51], and the Transformer-based encoders, including ViT [24] and Swin-T [64]. Table V shows the experimental results of the instruction encoders on the RSITMD dataset. The experimental results show that using ResNet-50 pre-trained on the AID dataset and RESISC45 dataset as the instruction encoder is more effective, performing the best at 38.24 and 38.28. We find no significant increase on retrieval performance by looking at the pre-trained ViT and Swin-T on the AID dataset as instruction encoders. By comparing VGG-16 and ResNet-50 pre-trained on ImageNet, we find that the pre-training based on remote sensing scene classification has a more obvious instructive role in guiding the vision representation. The above experimental results show that using prior knowledge based on remote sensing scene recognition can enhance vision representation and, thus, retrieval performance.
@@ V-G Instruction Strategy Analysis
VI CONCLUSION
This paper proposes the PIR framework, a new paradigm that leverages prior knowledge from the remote sensing data domain for adaptive learning of vision and text representations. We first propose the PIR-ITR for remote sensing image-text retrieval. Notably, Spatial-PAE and Temporal-PAE are applied to perform long-range dependency modelling to enhance key feature representation. In vision representation, we utilize prior knowledge from remote sensing scene recognition to instruct unbiased vision representation and reduce semantic noise’s impact. In text representation, Temporal-PAE activates the current time step based on the previous time step to enhance semantic representation capability. Finally, a cluster-wise attribution loss is used as an inter-class constraint to reduce semantic confusion zones in the common subspace. Building on this foundation, we propose PIR-CLIP to address semantic noise in remote sensing vision-language representations and enhance open-domain retrieval performance further. Comprehensive experiments are designed to validate the superiority and effectiveness of our approach on two benchmark datasets, RSICD and RSITMD.